What is the market price of a U.S. Treasury bond that has a
coupon rate of 9%, a face value of $1,000 and matures exactly
10 years from today if the required yield to maturity is 10%
compounded semiannually
1
Expert's answer
2020-07-21T18:40:54-0400

1.Present value of semiannual coupons:

"\frac{90}{0.1}"(1-"\frac{1 }{(1.05)20})" =$560.7


2.Present value of $1,000 principal:

"\frac{1000} {(1.05)20}" =$376.89

The price of the bond is the sum of the present values of coupons and principal:

Market Price = $560.7 +$376.89 =$ 937.59